# It’s Giving Major Satisfaction: Why Fairness Matters for Developers

### v1

@@ IV Survey Design
IV-A Measuring Perception of Fairness
In order to explore the relationship between fairness and job satisfaction among software practitioners, it is essential to first measure fairness effectively. Fairness, in this context, is understood as a personal judgment or evaluation of organizational activities, particularly regarding whether these activities are perceived as moral and praiseworthy [36]. To accurately assess this experience of fairness, we rely on the framework developed by Colquitt in 2001 [5].
Our survey design draws inspiration from the survey questions developed by Colquitt et al. [5], which consists of 20 questions. These questions assess various four dimensions of fairness, namely distributive, procedural, interpersonal, and informational fairness. Colquitt’s measure of fairness was developed through careful research, creating survey items based on key studies in the field of organizational justice, which examines fairness in workplaces. These items were tested in different settings, such as in a university classroom and a manufacturing company, to ensure they worked well across various environments [5].
By aligning with Colquitt’s well-established framework, we ensure that our approach is grounded in a reliable foundation. The 20 items used in our survey, as developed by Colquitt [5], are listed in Table I.
Table I: Questions to Measure Perception of Fairness [5]
@@ IV-A Measuring Perception of Fairness
IV-B Measuring Job Satisfaction
Given the connection between perceptions of fairness and job satisfaction across various fields, we inquire if the observed link between fairness and job satisfaction apply in the context of SE. To develop a more tailored approach to measuring job satisfaction in the SE field, we turn to the SE literature. Storey et al. [12] conducted a comprehensive survey specifically within the field of SE, aiming to identify the factors that most significantly impact job satisfaction among software engineers. Developed through two main iterations and rigorously piloted, the survey incorporated insights from a literature review, onsite observations, and internal survey comments. The final version, which gathered a total of 640 responses, identified 44 factors influencing job satisfaction in SE.
In designing our survey, we draw inspiration from the 44 job satisfaction factors and 1 overall job satisfaction question used by Storey et al. [12]. However, to ensure that the factors we examine are related to perceptions of fairness, we align our selection with organizational justice literature.
For example, one of the 44 factors is satisfaction with one’s manager. We intuitively believe that a manager’s behavior and leadership style play a significant role in how fairness is perceived by employees. For instance, during performance reviews, the way a manager communicates feedback, acknowledges contributions, and makes decisions about promotions or raises can significantly impact whether employees feel they are being treated fairly. Organizational justice literature supports this intuition, highlighting that a manager’s leadership style is a crucial factor in shaping employees’ perceptions of fairness within an organization [37]. Therefore, we believe that satisfaction with one’s manager is closely related to perceptions of fairness.
On the other hand, factor such as autonomy, which refers to an individual’s need for independence [16], may not directly relate to perceptions of fairness. Autonomy is more concerned with the freedom to carry out tasks and the ability to make decisions independently [16]. However, autonomy is primarily about personal control and is less about how fairly an organization’s policies, processes, and decisions are perceived by its employees. As such, autonomy does not align with the fairness constructs that are the focus of our study.
Consequently, we refined our scope from the original 44 factors to 15 that are more closely aligned with fairness, supported by empirical evidence. The selected factors, presented in Table II, allow us to analyze job satisfaction in a way that is directly linked to perceptions of fairness.
Table II: Key Satisfaction Items and Their Rationale in the Context of Organizational Fairness
@@ IV-B Measuring Job Satisfaction
Manager The authenticity of a manager’s leadership style plays a crucial factor in shaping employees’ perceptions of fairness within an organization [37].
@@
Feedback When employees perceive high levels of interactive fairness, the positive impact of developmental feedback on satisfaction and engagement will be amplified [38].
@@
Appreciation Satisfaction with appreciation shown for one’s work is closely linked to perceptions of fairness, which in turn may significantly impact an employee’s job satisfaction and performance [39].
@@
Priorities Employees’ perceptions of fairness in the organization can influence how they prioritize and engage with their work tasks [40].
@@
Organization Culture Procedural and distributive fairness can be strong predictors of satisfaction with work schedules [41].
@@
Team Culture When team members perceive the distribution of equity as fair, it may lead to positive interaction spirals within the team [42].
@@
Team Collaboration Higher perceptions of procedural and interactional fairness within an organization may be associated with better satisfaction with teamwork [43].
@@
Productivity Enhancing organizational fairness can increase staff productivity, recommending that administrators focus on improving components of organizational fairness [44].
@@
Ability In an environment where employees perceive high levels of fairness, they are more likely to take initiative and manage themselves effectively [45].
@@
Salary Larger financial rewards for high performers increase perceptions of distributive and procedural fairness, which in turn foster work motivation [46].
@@
Benefits Employees’ views on whether their compensation is fair are influenced not only by the amount they receive but also by how they perceive the process of determining that compensation [47].
@@
Rewards Distributive fairness related to pay, benefits, and rewards is significantly linked to job satisfaction [48].
@@
Work-Life Balance When decision-making processes are perceived as fair, employees are more likely to be satisfied with their work-life balance and less likely to consider leaving the organization [49].
@@
Promotion Employees’ perceptions of fairness in performance evaluations and promotion decisions are critical to their sense of value and belonging within an organization [50].
@@
Job Security When employees perceive fair treatment in the organization, they tend to feel more secure in their jobs [51].
IV-C Measuring Fairness Experiences in SE Contexts
@@
Our study measures perceptions of fairness using survey questions developed by Colquitt [5], grounded in the organizational justice literature’s theoretical framework for understanding fairness in the workplace. However, it is also important to understand how these fairness dimensions manifest in real-world scenarios, particularly within the context of SE. Exploring concrete examples of fairness issues in specific contexts can reveal where and how fairness is challenged, guiding the development of targeted strategies to improve job satisfaction.
@@
A study by Sesari et al. [4] identifies ten distinct contexts where fairness is a concern within the SE field: income, treatment, allocation of work, recruitment, working hours, demand, evaluation, authorship, policy, and customer/client relations. These contexts represent primary areas where fairness challenges are most likely to arise, offering valuable insights into where software engineers perceive fairness issues in their daily professional experiences. To investigate these contexts further, we asked survey respondents to rate the extent to which they have experienced fairness in each of the ten contexts identified by Sesari et al. [4]. This approach allows us to systematically measure the extent to which fairness perceptions are experienced in various SE scenarios.
IV-D Collecting Demographic Information
@@
Gender: Male, Female. Given the historical gender biases in the SE field, particularly in areas like hiring and promotion [53], we anticipate that gender may play a significant role in shaping job satisfaction.
@@
Age: Under 40, 40 or older. Older employees often report higher satisfaction due to increased experience, stability, and potentially more favorable job conditions [14]. This suggests that younger employees might face different challenges which could influence their job satisfaction differently.
@@
Ethnicity: American Indian/Alaska Native, Asian, Black, Hispanic, Mixed, Native Hawaiian/Pacific Islander, White, Other. Minority groups might face challenges such as discrimination or lack of representation, which can negatively impact their work experience [3, 54]. This could lead to varying levels of job satisfaction across different ethnicities, with minority groups potentially reporting lower satisfaction.
@@
Ethnicity Representation: Well-represented, Under-represented, in a Balanced/Diverse Team. Individuals whose ethnicity is under-represented in their groups may feel marginalized or distinct from their peers [55]. Those in well-represented or balanced teams may experience higher job satisfaction, while those in under-represented groups might report lower satisfaction due to feelings of isolation or discrimination.
@@
Work Experience: $\leq 1$ year, $1-3$ years, $3-5$ years, $5-7$ years, $>7$ years. More experienced individuals typically benefit from increased competence, more significant project roles, and stronger workplace relationships [56]. This could imply that those with less experience might face dissatisfaction, potentially due to less job security or fewer opportunities.
@@
Role: DevOps Engineer, Product Manager, Software Architect, Software Developer/Engineer, Software Tester/QA Engineer, Other. Roles that involve more autonomy and collaboration, such as Product Management, might contribute to higher job satisfaction [57], whereas roles with less autonomy or more routine tasks may see lower satisfaction levels.
@@
Role Nature: Management, Employee, Other. Managers might focus on the fairness of their decisions and policies, while employees are more likely to assess fairness in terms of how it affects them personally [58, 59]. This suggests that job satisfaction might vary between managers and employees.
@@
Work Limitation: Does a physical, mental, or emotional problem keep you from working? (yes/no) Are you limited in the kind or amount of work you can do because of a physical, mental, or emotional problem? (yes/no). Individuals with work limitations often perceive lower fairness, particularly in areas such as pay and workplace accommodations [60]. This might suggest that those with work limitations report lower job satisfaction compared to those without such limitations. The questions used in the survey are based on National Health Interview Survey (NHIS) [61].
IV-E Respondent Recruitment
@@
V Research Methods
In this section, we detail the methodologies employed to answer the research questions posed in this study. The scripts for all steps are provided in the supplementary material [62].
V-A Pre-processing
@@
The shortest response time was 0.6 minutes, which we deemed too short to yield meaningful data given the expected 10-minute duration. Research supports that very quick survey completions can indicate less attentive responses [63]. On reviewing the 0.6-minute response, we found missing answers in key areas, reinforcing concerns about data quality for very fast responses. We also examined responses with durations beyond the third quartile (Q3), ranging from 17 minutes to $\sim$6 days. The extended response times might be due to allowing respondents to return to the survey later, as their progress remains saved even if the tab is left open. Therefore, we did not find it necessary to exclude these, as longer completion times do not necessarily correlate with lower data quality.
@@
To do this, we applied Multiple Imputation by Chained Equations (MICE) using Predictive Mean Matching (PMM). This method fills in missing data by matching them with similar observed values, ensuring completeness and consistency in our dataset. MICE has been demonstrated to be versatile and robust in various research domains [65, 66, 67, 68, 69].
To ensure the reliability of our survey, we also conducted an Omega reliability test [70], which is well-suited for multidimensional surveys. Omega reliability test is based on a confirmatory factor analysis model that recognizes the possibility of multiple factors or dimensions within a survey [70]. In our survey, there are items representing job satisfaction, items representing fairness, and items representing fairness contexts. Given the multidimensional nature of these items, Omega reliability test was the appropriate choice to capture the varying contributions of these items across different dimensions.
@@ V-B Summarizing Responses
For the Likert responses to the 20-item fairness scale, analyzing individual items allows us to identify specific aspects of fairness that may need prioritization. However, to understand general trends across the four fairness dimensions, Enoksen [71] recommended aggregating the Likert responses by calculating the mean for each dimension. For example, we calculated the mean of the four items related to distributive fairness to create an aggregated descriptive statistic for this dimension.
To examine demographic differences in the three perceptions, we used the Mann-Whitney U [72] and Kruskal-Wallis [73] tests, along with wilcoxon_effsize [74] in R for effect size. We focused on moderate ($0.30-<0.5$) and large effect sizes ($>=0.5$) to highlight meaningful differences [74].
The Mann-Whitney U test compares responses between two groups [72], such as those with and without work limitations. For example, if the test yielded a p-value less than 0.05 and an effect size of 0.4 for satisfaction with the manager, it would indicate a moderate difference between the two groups. We then referred to the bar plot visualizations (see our supplementary material [62]) to explore these differences, such as the percentage of respondents with and without work limitations reacting negatively to satisfaction with the manager.
The Kruskal-Wallis test identifies statistically significant differences among three or more independent groups, such as respondents with varying levels of work experience [73]. When significant differences were detected, we conducted post hoc Dunn-Bonferroni tests to pinpoint specific group differences while controlling for Type I errors. For example, if the Kruskal-Wallis test indicated a p-value less than 0.05 for satisfaction with benefits across different work experience levels, and the post hoc test found a significant difference between respondents with less than 1 year and more than 7 years of experience with an effect size of 0.6, it would suggest a large effect. Similar to the Mann-Whitney U test, we then reviewed visualizations to examine these differences, such as the percentage of respondents with less than 1 year of experience versus those with more than 7 years reacting to satisfaction with benefits.
@@
V-C Grouping Job Satisfaction Factors

### v2

@@ 4. Survey Design
4.1. Measuring Perception of Fairness
In order to explore the relationship between fairness and job satisfaction among software practitioners, it is essential to first measure fairness effectively. Fairness, in this context, is understood as a personal judgment or evaluation of organizational activities, particularly regarding whether these activities are perceived as moral and praiseworthy (Goldman and Cropanzano, 2015). To accurately assess this experience of fairness, we rely on the framework developed by Colquitt in 2001 (Colquitt, 2001).
Our survey design draws inspiration from the survey questions developed by Colquitt et al. (Colquitt, 2001), which consists of 20 questions. These questions assess various four dimensions of fairness, namely distributive, procedural, interpersonal, and informational fairness. Colquitt’s measure of fairness was developed through careful research, creating survey items based on key studies in the field of organizational justice, which examines fairness in workplaces. These items were tested in different settings, such as in a university classroom and a manufacturing company, to ensure they worked well across various environments (Colquitt, 2001).
By aligning with Colquitt’s well-established framework, we ensure that our approach is grounded in a reliable foundation. The 20 items used in our survey, as developed by Colquitt (Colquitt, 2001), are listed in Table 1.
Table 1. Questions to Measure Perception of Fairness (Colquitt, 2001)
@@ 4.1. Measuring Perception of Fairness
4.2. Measuring Job Satisfaction
Given the connection between perceptions of fairness and job satisfaction across various fields, we inquire if the observed link between fairness and job satisfaction apply in the context of SE. To develop a more tailored approach to measuring job satisfaction in the SE field, we turn to the SE literature. Storey et al. (Storey et al., 2021) conducted a comprehensive survey specifically within the field of SE, aiming to identify the factors that most significantly impact job satisfaction among software engineers. Developed through two main iterations and rigorously piloted, the survey incorporated insights from a literature review, onsite observations, and internal survey comments. The final version, which gathered a total of 640 responses, identified 44 factors influencing job satisfaction in SE.
In designing our survey, we draw inspiration from the 44 job satisfaction factors and 1 overall job satisfaction question used by Storey et al. (Storey et al., 2021). However, to ensure that the factors we examine are related to perceptions of fairness, we align our selection with organizational justice literature.
For example, one of the 44 factors is satisfaction with one’s manager. We intuitively believe that a manager’s behavior and leadership style play a significant role in how fairness is perceived by employees. For instance, during performance reviews, the way a manager communicates feedback, acknowledges contributions, and makes decisions about promotions or raises can significantly impact whether employees feel they are being treated fairly. Organizational justice literature supports this intuition, highlighting that a manager’s leadership style is a crucial factor in shaping employees’ perceptions of fairness within an organization (Kiersch and Byrne, 2015). Therefore, we believe that satisfaction with one’s manager is closely related to perceptions of fairness.
On the other hand, factor such as autonomy, which refers to an individual’s need for independence (Beecham et al., 2008), may not directly relate to perceptions of fairness. Autonomy is more concerned with the freedom to carry out tasks and the ability to make decisions independently (Beecham et al., 2008). However, autonomy is primarily about personal control and is less about how fairly an organization’s policies, processes, and decisions are perceived by its employees. As such, autonomy does not align with the fairness constructs that are the focus of our study.
Consequently, we refined our scope from the original 44 factors to 15 that are more closely aligned with fairness, supported by empirical evidence. The selected factors, presented in Table 2, allow us to analyze job satisfaction in a way that is directly linked to perceptions of fairness.
Table 2. Key Satisfaction Items and Their Rationale in the Context of Organizational Fairness
@@ 4.2. Measuring Job Satisfaction
Manager The authenticity of a manager’s leadership style plays a crucial factor in shaping employees’ perceptions of fairness within an organization (Kiersch and Byrne, 2015).
@@
Feedback When employees perceive high levels of interactive fairness, the positive impact of developmental feedback on satisfaction and engagement will be amplified (Zhang and Zhu, 2019).
@@
Appreciation Satisfaction with appreciation shown for one’s work is closely linked to perceptions of fairness, which in turn may significantly impact an employee’s job satisfaction and performance (Shonubi et al., 2016).
@@
Priorities Employees’ perceptions of fairness in the organization can influence how they prioritize and engage with their work tasks (Cropanzano et al., [n. d.]).
@@
Organization Culture Procedural and distributive fairness can be strong predictors of satisfaction with work schedules (Zeidler et al., 2021).
@@
Team Culture When team members perceive the distribution of equity as fair, it may lead to positive interaction spirals within the team (Breugst et al., 2015).
@@
Team Collaboration Higher perceptions of procedural and interactional fairness within an organization may be associated with better satisfaction with teamwork (Dayan and Benedetto, 2008).
@@
Productivity Enhancing organizational fairness can increase staff productivity, recommending that administrators focus on improving components of organizational fairness (Hosseini et al., 2016).
@@
Ability In an environment where employees perceive high levels of fairness, they are more likely to take initiative and manage themselves effectively (Park et al., 2016).
@@
Salary Larger financial rewards for high performers increase perceptions of distributive and procedural fairness, which in turn foster work motivation (Sebastian, 2016).
@@
Benefits Employees’ views on whether their compensation is fair are influenced not only by the amount they receive but also by how they perceive the process of determining that compensation (Wu and Wang, 2008).
@@
Rewards Distributive fairness related to pay, benefits, and rewards is significantly linked to job satisfaction (Haar and Spell, 2009).
@@
Work-Life Balance When decision-making processes are perceived as fair, employees are more likely to be satisfied with their work-life balance and less likely to consider leaving the organization (Cao et al., 2013).
@@
Promotion Employees’ perceptions of fairness in performance evaluations and promotion decisions are critical to their sense of value and belonging within an organization (Turhan et al., 2016).
@@
Job Security When employees perceive fair treatment in the organization, they tend to feel more secure in their jobs (Teimouri et al., 2016).
4.3. Measuring Fairness Experiences in SE Contexts
@@
Our study measures perceptions of fairness using survey questions developed by Colquitt (Colquitt, 2001), grounded in the organizational justice literature’s theoretical framework for understanding fairness in the workplace. However, it is also important to understand how these fairness dimensions manifest in real-world scenarios, particularly within the context of SE. Exploring concrete examples of fairness issues in specific contexts can reveal where and how fairness is challenged, guiding the development of targeted strategies to improve job satisfaction.
@@
A study by Sesari et al. (Sesari et al., 2024) identifies ten distinct contexts where fairness is a concern within the SE field: income, treatment, allocation of work, recruitment, working hours, demand, evaluation, authorship, policy, and customer/client relations. These contexts represent primary areas where fairness challenges are most likely to arise, offering valuable insights into where software engineers perceive fairness issues in their daily professional experiences. To investigate these contexts further, we asked survey respondents to rate the extent to which they have experienced fairness in each of the ten contexts identified by Sesari et al. (Sesari et al., 2024). This approach allows us to systematically measure the extent to which fairness perceptions are experienced in various SE scenarios.
4.4. Collecting Demographic Information
@@
Gender: Male, Female. Given the historical gender biases in the SE field, particularly in areas like hiring and promotion (Wang and Redmiles, 2019), we anticipate that gender may play a significant role in shaping job satisfaction.
@@
Age: Under 40, 40 or older. Older employees often report higher satisfaction due to increased experience, stability, and potentially more favorable job conditions (Baltes et al., 2020). This suggests that younger employees might face different challenges which could influence their job satisfaction differently.
@@
Ethnicity: American Indian/Alaska Native, Asian, Black, Hispanic, Mixed, Native Hawaiian/Pacific Islander, White, Other. Minority groups might face challenges such as discrimination or lack of representation, which can negatively impact their work experience (Nadri et al., 2021; Shameer et al., 2023). This could lead to varying levels of job satisfaction across different ethnicities, with minority groups potentially reporting lower satisfaction.
@@
Ethnicity Representation: Well-represented, Under-represented, in a Balanced/Diverse Team. Individuals whose ethnicity is under-represented in their groups may feel marginalized or distinct from their peers (Choi, 2016). Those in well-represented or balanced teams may experience higher job satisfaction, while those in under-represented groups might report lower satisfaction due to feelings of isolation or discrimination.
@@
Work Experience: $\leq 1$ year, $1-3$ years, $3-5$ years, $5-7$ years, $>7$ years. More experienced individuals typically benefit from increased competence, more significant project roles, and stronger workplace relationships (Tarasov, 2019). This could imply that those with less experience might face dissatisfaction, potentially due to less job security or fewer opportunities.
@@
Role: DevOps Engineer, Product Manager, Software Architect, Software Developer/Engineer, Software Tester/QA Engineer, Other. Roles that involve more autonomy and collaboration, such as Product Management, might contribute to higher job satisfaction (Yilmaz et al., 2012), whereas roles with less autonomy or more routine tasks may see lower satisfaction levels.
@@
Role Nature: Management, Employee, Other. Managers might focus on the fairness of their decisions and policies, while employees are more likely to assess fairness in terms of how it affects them personally (Gilliland et al., 2008; Barclay et al., 2005). This suggests that job satisfaction might vary between managers and employees.
@@
Work Limitation: Does a physical, mental, or emotional problem keep you from working? (yes/no) Are you limited in the kind or amount of work you can do because of a physical, mental, or emotional problem? (yes/no). Individuals with work limitations often perceive lower fairness, particularly in areas such as pay and workplace accommodations (Milner et al., 2015). This might suggest that those with work limitations report lower job satisfaction compared to those without such limitations. The questions used in the survey are based on National Health Interview Survey (NHIS) (Altman, 2014).
4.5. Respondent Recruitment
@@
5. Research Methods
In this section, we detail the methodologies employed to answer the research questions posed in this study. The scripts for all steps are provided in the supplementary material (fig, 2024).
5.1. Pre-processing
@@
The shortest response time was 0.6 minutes, which we deemed too short to yield meaningful data given the expected 10-minute duration. Research supports that very quick survey completions can indicate less attentive responses (Wood et al., 2017). On reviewing the 0.6-minute response, we found missing answers in key areas, reinforcing concerns about data quality for very fast responses. We also examined responses with durations beyond the third quartile (Q3), ranging from 17 minutes to $\sim$6 days. The extended response times might be due to allowing respondents to return to the survey later, as their progress remains saved even if the tab is left open. Therefore, we did not find it necessary to exclude these, as longer completion times do not necessarily correlate with lower data quality.
@@
To do this, we applied Multiple Imputation by Chained Equations (MICE) using Predictive Mean Matching (PMM). This method fills in missing data by matching them with similar observed values, ensuring completeness and consistency in our dataset. MICE has been demonstrated to be versatile and robust in various research domains (White et al., 2011; Chhabra et al., 2017; Wulff and Jeppesen, 2017; Azur et al., 2011; Beesley et al., 2021).
To ensure the reliability of our survey, we also conducted an Omega reliability test (McDonald, 1970), which is well-suited for multidimensional surveys. Omega reliability test is based on a confirmatory factor analysis model that recognizes the possibility of multiple factors or dimensions within a survey (McDonald, 1970). In our survey, there are items representing job satisfaction, items representing fairness, and items representing fairness contexts. Given the multidimensional nature of these items, Omega reliability test was the appropriate choice to capture the varying contributions of these items across different dimensions.
@@ 5.2. Summarizing Responses
For the Likert responses to the 20-item fairness scale, analyzing individual items allows us to identify specific aspects of fairness that may need prioritization. However, to understand general trends across the four fairness dimensions, Enoksen (Enoksen, 2015) recommended aggregating the Likert responses by calculating the mean for each dimension. For example, we calculated the mean of the four items related to distributive fairness to create an aggregated descriptive statistic for this dimension.
To examine demographic differences in the three perceptions, we used the Mann-Whitney U (MacFarland and Yates, 2016) and Kruskal-Wallis (Kruskal and Wallis, 1952) tests, along with wilcoxon_effsize (Tomczak and Tomczak, 2014) in R for effect size. We focused on moderate ($0.30-<0.5$) and large effect sizes ($>=0.5$) to highlight meaningful differences (Tomczak and Tomczak, 2014).
The Mann-Whitney U test compares responses between two groups (MacFarland and Yates, 2016), such as those with and without work limitations. For example, if the test yielded a p-value less than 0.05 and an effect size of 0.4 for satisfaction with the manager, it would indicate a moderate difference between the two groups. We then referred to the bar plot visualizations (see our supplementary material (fig, 2024)) to explore these differences, such as the percentage of respondents with and without work limitations reacting negatively to satisfaction with the manager.
The Kruskal-Wallis test identifies statistically significant differences among three or more independent groups, such as respondents with varying levels of work experience (Kruskal and Wallis, 1952). When significant differences were detected, we conducted post hoc Dunn-Bonferroni tests to pinpoint specific group differences while controlling for Type I errors. For example, if the Kruskal-Wallis test indicated a p-value less than 0.05 for satisfaction with benefits across different work experience levels, and the post hoc test found a significant difference between respondents with less than 1 year and more than 7 years of experience with an effect size of 0.6, it would suggest a large effect. Similar to the Mann-Whitney U test, we then reviewed visualizations to examine these differences, such as the percentage of respondents with less than 1 year of experience versus those with more than 7 years reacting to satisfaction with benefits.
@@
5.3. Grouping Job Satisfaction Factors